Please solve
2X-9=5x-7

Answer by checkley77(12844)

2X-9=5x-7
2x-5x=-7+9
-3x=2
x=2/-3
x=-2/3 answer.
Proof:
2*-2/3-9=5*-2/3-7
-4/3-9=-10/3-7
10/3-4/3=9-7
6/3=2
2=2